# Flexible and Comprehensive Evaluation for LLMs

### v1

@@ Abstract
Evaluation is pivotal for honing Large Language Models (LLMs), pinpointing their capabilities and guiding enhancements. The rapid development of LLMs calls for a lightweight and easy-to-use framework for swift evaluation deployment. However, due to the various implementation details to consider, developing a comprehensive evaluation platform is never easy. Existing platforms are often complex and poorly modularized, hindering seamless incorporation into researcher’s workflows. This paper introduces UltraEval, a user-friendly evaluation framework characterized by lightweight, comprehensiveness, modularity, and efficiency. We identify and reimplement three core components of model evaluation (models, data, and metrics). The resulting composability allows for the free combination of different models, tasks, prompts, and metrics within a unified evaluation workflow. Additionally, UltraEval supports diverse models owing to a unified HTTP service and provides sufficient inference acceleration. UltraEval is now available for researchers publicly  [^1] Website is at https://github.com/OpenBMB/UltraEval and a demo video is at https://youtu.be/C0O6BVzNAS8..
@@
[…]
@@ 1 Introduction
This advancement is significantly bridging the gap between the realization of Artificial General Intelligence (AGI)Bubeck et al. (2023) and the current state. Nevertheless, the expansion of model parameters and training datasets engenders increasing uncertainties and emergent capabilities, posing potential risks to humanity and challenges to stable training models Chang et al. (2023); Bommasani et al. (2021); Wei et al. (2022a). Consequently, it is imperative to continuously and meticulously evaluate the evolving capabilities of LLMs throughout their development to ensure their responsible and beneficial application.
@@
Currently, some evaluation frameworks covering the entire pipeline from model deployment to model evaluation have been proposed, and are predominantly divided into two types: conversational websites, exemplified by platforms like Chatbot Arena  [^2] https://chat.lmsys.org/, and open-source evaluation tools, such as lm-evaluation-harness  [^3] https://github.com/EleutherAI/lm-evaluation-harness. The former effectively assesses the conversational abilities of a model but lacks comprehensive task coverage and transparency in the evaluation process. The latter open-source frameworks face challenges such as incomplete task coverage, complex code structure, heavy implementations, high difficulty to use, and tightly coupled functionalities. These issues hinder both convenient and comprehensive assessments.
@@
Table 1: Comparison of Evaluation Frameworks. Bytes: the total number of bytes of Python and Jupyter Notebook code within each framework’s GitHub repository, Acceleration: tools or methods employed to expedite model inference, Model Types: Supported Models for Evaluation, GPT: GPT series models, LLaMA: LLaMA series models
In this paper, we identify three core components that form an evaluation process: models (or systems), task data, and metrics (i.e., evaluation methods), as illustrated in Figure 1. Rethinking the implementations in these three aspects would benefit the construction of a lightweight and easy-to-use evaluation framework, which covers mainstream tasks and complete evaluation pipeline, and can be easily expanded according to user customization. To this end, we introduce UltraEval, a lightweight and user-friendly open-source framework for LLMs evaluation. It stands out for its modular and scalable design, enabling thorough assessment of model capabilities. As illustrated in Figure 2, we segment the evaluation pipeline into three main modules: Data, Model, and Metrics, each operating independently and interacting through data exchange.
@@ 3.1 Data preparation
Data Preprocessing. Initially, we collect commonly used benchmarks for evaluating LLMs, such as MMLUHendrycks et al. (2020), GSM8KCobbe et al. (2021), and HellaswagZellers et al. (2019), covering multiple dimensions of capabilities. Currently, we have 59 benchmarks, listed in Table 2, and we plan to continually expand our collection of benchmarks.
@@
Prompt Templates. Prompts are used to guide AI models to generate specific outputs, and UltraEval supports prompt engineering White et al. (2023), including few-shot and Chain of Thought (CoT) Wei et al. (2022b), to enhance the model’s accuracy in reasoning. The sensitivity of LLMs to prompts Zhu et al. (2023) and the variability of prompts across different tasks often make it challenging for researchers to replicate results from papers, hindering research progress. UltraEval addresses this issue by providing customized, stable prompt templates for each task to facilitate result alignment. Figure 4 showcases an example of a prompt template for MMLU, demonstrating a rigorous process for forming the final prompt input.
3.2 Model delpoyment
@@
Http Service. In traditional evaluations, model deployment is closely integrated with task assessment, requiring models to be redeployed for each new task, which can lead to unnecessary consumption of time and hardware resources. To address this, we deploy models as HTTP services with Flask, facilitating modularization and efficient resource use. This approach offers several advantages:
@@
Multi-GPU Acceleration. We leverage the Gunicorn web server with Flask to deploy models via web endpoints, achieving a flexible and decoupled architecture for model deployment and evaluation. This setup allows for dynamic GPU acceleration, where the Gunicorn server, configured with environment-specific parameters, manages multiple worker processes. Each process, handling a slice of the available GPUs, executes inference tasks in parallel, significantly improving computational efficiency and throughput. A highlight of UltraEval’s performance is its ability to utilize 4 A800 GPUs to evaluate a test set of 41k data points in under 1.5 hours, showcasing remarkable efficiency [^16] The model used in this experiment is Llama2-7b, and the benchmarks includes BBH, MMLU, C-Eval, CMMLU, HumanEval, MBPP, GSM8K and MATH. In total, they consist of 40938 data points..
@@ 3.3 Evaluation methods
Post-Processing. Model outputs, influenced by task characteristics, prompt templates, and the model’s performance, often contain extraneous information beyond the answers neededPark et al. (2024). As shown in Figure 7, when ChatGPT responds to HumanEval questions, the response may include both code and additional textual descriptions, complicating automatic evaluation. To more accurately assess model, it is necessary to post-process the model’s outputs to extract the most crucial answers.
@@
ARC-C 45.9 43.17 45.9 47.44 55.5 50.77
HellaSwag 77.2 75.64 80.7 79.13 81.3 80.43
BBH 32.6 32.8 39.4 39.17 $\approx$38 40.35
MATH 2.5 2.75 3.9 4.77 13.1 10.19
GSM8K 14.6 14.78 28.7 22.59 $52.1^{*}$ 31.92
HumanEval 12.8 12.8 18.3 17.07 30.5 26.83
MBPP 20.8 20.84 30.6 29.04 47.5 47.31
MMLU 45.3 45.14 54.8 55.18 60.1 63.05
Table 3: Evaluation results on mainstream benchmarks. $*{}^{*}$ The replicated result for GSM8K is 35.4 in Gemma paperTeam et al. (2024), which is close to our result.
@@ 4 Evaluation
UltraEval aims to provide a lightweight, comprehensive, and user-friendly evaluation framework to support research. As illustrated in Figure 5, UltraEval’s modular design effectively combines various models, tasks and metrics for evaluation. Utilizing UltraEval, we evaluate models from the LLaMA2 seriesTouvron et al. (2023) and MistralJiang et al. (2023) on these widely-used benchmarks. As indicated in the Table 3, our replicated results align with those reported in the literature, underscoring our framework’s reliability.
Furthermore, UltraEval supports innovative research efforts, research on predictable scalingHu et al. (2023), OlympiadBenchHe et al. (2024) and model training, such as with MiniCPMmin (2024).
5 Conclusion and Future Work
We introduce UltraEval, a lightweight, user-friendly, and comprehensive framework for model evaluation. UltraEval establishes a unified structure with well-defined modules and flexible interactions, aiding researchers and developers in efficiently deploying evaluation workflows. Moving forward, we plan to continuously integrate new technologies and features into UltraEval, extending beyond large language models to support multimodal, Retrieval-Augmented Generation (RAG), agents, and more, to advance the research on AGI. Additionally, we aim to expand our collection of representative benchmarks and also develop our own, exploring the capabilities and limits of large models.
@@ 5 Conclusion and Future Work
Presently, our approach primarily utilizes text domain evaluation. However, we are looking to expand the scope of UltraEval by integrating multi-modal and long-context evaluation datasets. This enhancement aims to facilitate more thorough and diverse assessments. Additionally, there is room for improvement in the visualization of our results. Future improvement will focus on enabling multi-dimensional visualization, thereby enriching the interpretability and depth of our evaluation results.

### v3

@@ Abstract
Evaluation is pivotal for refining Large Language Models (LLMs), pinpointing their capabilities, and guiding enhancements. The rapid development of LLMs calls for a lightweight and easy-to-use framework for swift evaluation deployment. However, considering various implementation details, developing a comprehensive evaluation platform is never easy. Existing platforms are often complex and poorly modularized, hindering seamless incorporation into research workflows. This paper introduces UltraEval, a user-friendly evaluation framework characterized by its lightweight nature, comprehensiveness, modularity, and efficiency. We identify and reimplement three core components of model evaluation (models, data, and metrics). The resulting composability allows for the free combination of different models, tasks, prompts, benchmarks, and metrics within a unified evaluation workflow. Additionally, UltraEval supports diverse models owing to a unified HTTP service and provides sufficient inference acceleration. UltraEval is now available for researchers publicly  [^1] The website of UltraEval is at https://github.com/OpenBMB/UltraEval and a demo video is at https://youtu.be/C0O6BVzNAS8..
@@
Chaoqun He1, Renjie Luo2222Work done during internship at ModelBest Inc., Shengding Hu1, Yuanqian Zhao3222Work done during internship at ModelBest Inc., Jie Zhou4 Hanghao Wu4, Jiajie Zhang5222Work done during internship at ModelBest Inc., Xu Han1∗,Zhiyuan Liu1thanks: Corresponding author:Xu Han (thu.hanxu13@gmail.com) and Zhiyuan Liu (liuzy@tsinghua.edu.cn), Maosong Sun1 1 Dept. of Comp. Sci. & Tech., Institute for AI, Tsinghua University, Beijing, China 2 Institute of Artificial Intelligence, Beihang University, China 3 Renmin University of China 4 ModelBest Inc. 5 Northeastern University, China {hcq21, hsd23}@mails.tsinghua.edu.cn, renjie.luo@outlook.com
@@ 1 Introduction
This advancement is significantly bridging the gap between the realization of the current state and Artificial General Intelligence (AGI)Bubeck et al. (2023). Nevertheless, the expansion of model parameters and training datasets engenders increasing uncertainties and emergent capabilities, posing potential risks to humanity and challenges to stable training models Chang et al. (2023); Bommasani et al. (2021); Wei et al. (2022a). Consequently, it is imperative to continuously and meticulously evaluate the evolving capabilities of LLMs throughout their development to ensure their responsible and beneficial applications.
@@
Currently, some evaluation frameworks covering the entire pipeline from model deployment to model evaluation are proposed, and predominantly divided into two types: conversational websites, exemplified by platforms like Chatbot Arena  [^2] https://chat.lmsys.org/, and open-source evaluation tools, such as lm-evaluation-harness  [^3] https://github.com/EleutherAI/lm-evaluation-harness. The former effectively assesses the conversational abilities of a model but lacks comprehensive task coverage and transparency in the evaluation process. The latter open-source frameworks face challenges such as incomplete task coverage, complex code structure, heavy implementations, high difficulty to use, and tightly coupled functionalities. These issues hinder both convenient and comprehensive assessments.
@@
Table 1: Comparison of Evaluation Frameworks. Bytes: the total bytes of Python and Jupyter Notebook code in each framework’s GitHub repository, Acceleration: tools or methods employed to expedite model inference, Model Types: Supported Models for Evaluation, GPT: GPT series models, LLaMA: LLaMA series models
In this paper, we identify three core components that form the evaluation process: models (or systems), task data, and metrics (i.e., evaluation methods), as illustrated in Figure 1. Rethinking the implementations in these three aspects would benefit the construction of a lightweight and easy-to-use evaluation framework, which covers mainstream tasks and complete evaluation pipeline, and can be easily expanded according to user customization. To this end, we introduce UltraEval, a lightweight and user-friendly open-source framework for LLMs evaluation. It stands out for its modular and scalable design, enabling thorough assessment of model capabilities. As illustrated in Figure 2, we segment the evaluation pipeline into three main modules: Data, Model, and Metrics, each operating independently and interacting through data exchange.
@@ 3.1 Data Preparation
Data Preprocessing. We collect commonly used benchmarks for evaluating LLMs, such as MMLUHendrycks et al. (2020), GSM8KCobbe et al. (2021), and HellaswagZellers et al. (2019), covering multiple dimensions of capabilities. Currently, we have 59 benchmarks, listed in Table 2, and we plan to continually expand our collection of benchmarks.
@@
Prompt Templates. Prompts are used to guide models to generate specific outputs, and UltraEval supports prompt engineering White et al. (2023), including few-shot and Chain of Thought (CoT) Wei et al. (2022b), to enhance the model’s accuracy. The sensitivity of LLMs to prompts Zhu et al. (2023) and the variability of prompts across different tasks often make it challenging for researchers to replicate results from papers, hindering research progress. UltraEval addresses this issue by providing customized, stable prompt templates for each task to facilitate result alignment. Figure 4 showcases an example of a prompt template for MMLU, demonstrating the rigorous process for forming the final prompt input.
3.2 Model Deployment
@@
Http Service. In traditional evaluations, model deployment is closely integrated with task assessment, requiring models to be redeployed for each new task, which can lead to unnecessary consumption of time and hardware resources. To address this, we deploy models as HTTP services with Flask, facilitating modularization and efficient resource use. This approach has several advantages:
@@
Multi-GPU Acceleration. We use the Gunicorn web server with Flask to deploy models via web endpoints, achieving a flexible and decoupled architecture for model deployment and evaluation. This setup allows for dynamic GPU acceleration, where the Gunicorn server, configured with environment-specific parameters, manages multiple worker processes. Each process, handling a slice of the available GPUs, executes inference tasks in parallel, significantly improving computational efficiency and throughput. A highlight of UltraEval’s performance is its ability to utilize 4 A800 GPUs to evaluate a test set of 41k data points in under 1.5 hours, showcasing remarkable efficiency [^16] The model used in this experiment is Llama2-7B, and the benchmarks includes BBH, MMLU, C-Eval, CMMLU, HumanEval, MBPP, GSM8K and MATH. In total, they consist of 40,938 data points..
@@ 3.3 Evaluation Methods
Post-Processing. Model outputs, influenced by task characteristics, prompt templates, and the model’s performance, often contain extraneous information beyond the answers neededPark et al. (2024). As shown in Figure 7, when ChatGPT responds to HumanEval questions, the response may include both code and additional textual descriptions, complicating automatic evaluation. To more accurately assess the model, it is necessary to post-process the model’s outputs to extract the most crucial answers.
@@
ARC-C 45.9 43.2 45.9 47.4 55.5 50.8
HellaSwag 77.2 75.6 80.7 79.1 81.3 80.4
BBH 32.6 32.8 39.4 39.2 $38.0^{*}$ 40.4
MATH 2.5 2.8 3.9 4.8 13.1 10.2
GSM8K 14.6 14.8 28.7 22.6 $52.1^{*}$ 31.9
HumanEval 12.8 12.8 18.3 17.1 30.5 26.8
MBPP 20.8 20.8 30.6 29.0 47.5 47.3
MMLU 45.3 45.1 54.8 55.2 60.1 63.1
Table 3: Evaluation results on mainstream benchmarks (%). The BBH score is not explicitly stated in the paperJiang et al. (2023), however, it is inferred to be 38.0 from the figures in the paper. The replicated result for GSM8K is 35.4 in Gemma paperTeam et al. (2024), which is close to our result.
@@ 4 Evaluation
UltraEval aims to provide a lightweight, comprehensive, and user-friendly evaluation framework to support research. As illustrated in Figure 5, UltraEval’s modular design effectively combines various models, tasks and metrics for evaluation. Using UltraEval, we evaluate models from the LLaMA2 seriesTouvron et al. (2023) and MistralJiang et al. (2023) on these widely-used benchmarks. As indicated in the Table 3, some reproduced results are higher, while others are lower, but within a certain margin of error, our reproduced results are consistent with the results reported in the papers, underscoring our framework’s reliability. The sources of error include hyperparameters (e.g., temperature, top-p) and hardware configurations. Since the evaluation details are not provided in the papers, verification is not possible. This highlights the importance of having an open and reproducible evaluation framework.
Furthermore, UltraEval supports innovative research efforts, research on predictable scalingHu et al. (2023), OlympiadBenchHe et al. (2024) and model training, such as with MiniCPMHu et al. (2024).
5 Discussion and Future Work
In this section, we discuss future work following this study. Specifically, we focus on addressing data contamination and supporting a broader range of evaluation scenarios.
Data contamination refers to the phenomenon that examples from the evaluation set are also found in the training data Li et al. (2024), causing inaccuracies in model evaluation. Common methods for contamination detection include n-gram overlap and embedding similarity search. However, none of these methods are perfect, making research on contamination detection still crucial. We will integrate these methods.
Supporting more evaluation scenarios, such as multimodal, long-text, and Retrieval-Augmented Generation (RAG), is crucial for meeting a broader range of evaluation needs. This will be an important direction for our future work.
6 Conclusion
We introduce UltraEval, a lightweight, user-friendly, and comprehensive framework for model evaluation. UltraEval establishes a unified structure with well-defined modules and flexible interactions, aiding researchers and developers in efficiently deploying evaluation workflows. Moving forward, we plan to continuously integrate new technologies and features into UltraEval, extending beyond large language models to support the evaluation of multimodal models, Retrieval-Augmented Generation (RAG), agents, and more, to advance the research on AGI. Additionally, we aim to expand our collection of representative benchmarks and also develop our own, exploring the capabilities and limits of large models.
Acknowledgements
This work is supported by the National Key R&D Program of China (No. 2022ZD0116312), Quan Cheng Laboratory (Grant No. QCLZD202301), the Postdoctoral Fellowship Program of CPSE (Grant No. GZB20230343), China Postdoctoral Science Foundation (Grant No. 2023M741945), and Institute Guo Qiang at Tsinghua University.
@@ Acknowledgements
Currently, our approach primarily utilizes text domain evaluation. However, we are looking to expand the scope of UltraEval by integrating multi-modal and long-context evaluation datasets. This enhancement aims to facilitate more thorough and diverse assessments. Additionally, there is room for improvement in the visualization of our results. Future improvement will focus on enabling multi-dimensional visualization, thereby enriching the interpretability and depth of our evaluation results.
@@ A.1 Post-process
A.2 UltraEval Usage
In the tutorial  [^17] https://github.com/OpenBMB/UltraEval/blob/main/docs/tutorials/en/ultraeval.md, we provide a detailed guide on UltraEval, including an introduction to its modules and instructions on how users can customize evaluations, such as adding their own new tasks and new models. This ensures that diverse evaluation needs are met.
A.3 Model as Judge
Currently, UltraEval supports ChatGPT as a substitute for human evaluation. However, since ChatGPT is a commercial model, users need to provide the relevant API key. Due to the modularity and flexibility of our framework, users can also use other closed-source models or highly performant open-source models as evaluators.
A.4 Multi-GPU Acceleration
Multi-GPU acceleration requires sufficient GPU resources. For larger models, such as Llama2-70B, deploying a single instance requires two A100 GPUs. Therefore, when using Multi-GPU acceleration, it is essential to consider both the model and the available GPU resources to ensure optimal configuration.